OpenStudy (anonymous):

Mr. Brown is creating examples of systems of equations. He wants to have a system of equations with infinite solutions that includes the equation 5x + 2y = 8. Which equation could Mr. Brown use to complete the system so that it has infinite solutions? 2x + 5y = 8 15x + 6y = 21 1.25x +0.5y = 2 6.5x + 3.5y = 9.5

OpenStudy (welshfella):

the correct equation would generate the same line on a graph and it would be a multiple of your equation for example x + y = 2 and 2x + 2y = 4 would have infinite solutions

OpenStudy (anonymous):

Yeah I don't know how to solve this problem, its confusing mee lol

OpenStudy (welshfella):

Hint: the equation could also be one where each coefficent of your equation is divided by the same number

OpenStudy (welshfella):

eg x + y = 2 and 0.5x + 0.5y = 1

OpenStudy (welshfella):

- where every term is divided by 2

OpenStudy (welshfella):

Try dividing your equation by 4 what do you get?